1 the oxidation state of the hydrogen atoms in CaH2 is

2 The oxidation number of the manganese atom in (NH4)2Mn04 I is

3 The oxidation number of phosphorus in KH2PO4 is

4 The oxidation number of the oxygen in O3 is

5 The oxidation number of p in H3PO2

6 Type of reagent is required to convert NO2- to N2 ?

7 what type of reagent is required to convert No2- to HNO2 ?

8 What is the total number of atoms contained in 2.00 of nickel?


1
Expert's answer
2021-05-31T05:58:04-0400

1)+1 × 2 = +2

2) -1 +(2×4) = +7

3) +5

4) (-1 + 1 + 0)/3 = 0

5) +1

6) Reducing agent

7) Oxidizing agent

8) One mole of any substance contains the same numbers of particles, which is known as Avogadro's constant

Particle=12.046×1023 molecules×1 atomparticle=12.0×1023 atoms≈12.0×1023 atoms

Hence, 2.00 moles of nickel contain 12.0×1023 Atoms